Find the Potential Energy of a charge with 99 μC with a potential difference of 15 V



1
Expert's answer
2022-03-20T18:50:38-0400

The Potential Energy of a charge 

"E_p=\frac{qV}{2}=\rm\frac{(99\times 10^{-6}\: C)\times(15\: V)}{2}=7.4\times 10^{-4}\: J"